マイクロメカニクス工学 2 単位 (選択)

Micromechanics  松尾繁樹 · 准教授/ 環境創生工学専攻エコシステム工学コース資源循環工学講座

【授業目的】マイクロ · ナノメートル空間領域の科学を理解するとともに，その ; 領域を取り扱う要素技術に関する知識を得ること． ; 【授業概要】マイクロメートルおよびナノメートル空間領域における精密科学計 ; 測手法について述べるとともに，マイクロマシンの要素技術であるマイクロ ; センサ，マイクロマシニング，マイクロアクチュエータの各技術についても ; 解説する．加えて，科学および技術に関する最新の話題にも随時言及し，議 ; 論する．本科目は，工業に関する科目である． ; 【授業形式】講義 ; 【キーワード】マイクロ，ナノ ; 【先行科目】 [先行科目] ; 【関連科目】 [関連科目] ; 【履修要件】 [要件] ; 【履修上の注意】 [注意] ; 【到達目標】マイクロ · ナノメートル空間を取り扱う基本的な知識を身につける ; こと． ; 【授業計画】 ; 1. マイクロスケールの物理と化学 ; 2. マイクロスケールの物理と化学 ; 3. マイクロスケールの物理と化学 ; 4. マイクロスケールの物理と化学 ; 5. マイクロスケールの計測技術 ; 6. マイクロスケールの計測技術 ; 7. マイクロスケールの計測技術 ; 8. マイクロマシニング技術 ; 9. マイクロマシニング技術 ; 10. マイクロマシニング技術 ; 11. マイクロセンサ ; 12. マイクロセンサ ; 13. マイクロセンサ ; 14. マイクロアクチュエータ ; 15. マイクロアクチュエータ ; 【成績評価基準】授業時間中に課すレポートで評価する． ; 【教科書】 [教科書] 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216887 【対象学生】 [対象学生] 【連絡先】 [連絡先]

Micromechanics 2 units (selection)

Shigeki Matsuo · Associate Professor / Resource Circulatory Engineering, Ecosystem Engineering, Earth and Life Environmental Engineering

Target⟩ To understand the science of micro-to-nanometer domain, and learn the

techniques to investigate such a domain. Outline⟩ To study the methods for precise measurement and manufacturing, in

addition, microsensors, micromachining, and microactuators. In addition, hot topics concerning science and technology will be referred and discussed at any time.

Style⟩ Lecture �������⟩ micro, nano Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ To obtain basic knowledges on the micro-to-nanometer domain

Schedule⟩

1. Physics and chemistry in microdomain

2. Physics and chemistry in microdomain

3. Physics and chemistry in microdomain

4. Physics and chemistry in microdomain

5. Measurements in microdomain

6. Measurements in microdomain

7. Measurements in microdomain

8. Micromachining

9. Micromachining

10. Micromachining

11. Microsensors

12. Microsensors

13. Microsensors

14. Microactuators

15. Microactuators Evaluation Criteria⟩ Assignments count 100%. Textbook⟩ [教科書] Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216887 Student⟩ [対象学生] Contact⟩ [連絡先]